OpenStudy (anonymous):

the lenght of a rectangle is 2 centimenters longer than its width. the perimeter is 16 centimenters. what are the lenght and width of the re

OpenStudy (shadowfiend):

Can you try to make this information into two equations?

OpenStudy (anonymous):

not really. I don't know where to begin

OpenStudy (shadowfiend):

Ok, so if the length is 2 cm longer than its width, we can say that: \[l = w + 2\text{ cm}\] Right? What does the perimeter tell us?

OpenStudy (anonymous):

I am still stuck

OpenStudy (shadowfiend):

Well, what is the perimeter of a rectangle?

OpenStudy (anonymous):

l+w=16 ?

OpenStudy (shadowfiend):

Exactly. (16 cm, in this case.)

OpenStudy (anonymous):

ok. then what do I do?

OpenStudy (anonymous):

w+2cm+w=16cm?

OpenStudy (anonymous):

is that right so far

OpenStudy (shadowfiend):

Sorry, yes, that's right.